OpenStudy (anonymous):

Find the equation of a 3rd degree polynomial with zeros 2 and 3i. And f(1)=3

OpenStudy (anonymous):

start with \[a(x-2)(x-3i)(x+3i)\] and then multiply out. it is not too hard since \[(x-3i)(x+3i)=x^2+9\]

OpenStudy (anonymous):

you should get \[a(x^3-2 x^2+9 x-18)\] and you know if x = 1, then the result is 3, so replace x by 1 and set the answer = 3 and solve for a

OpenStudy (anonymous):

i get \[-10a=3\] so \[a=-\frac{3}{10}\] but you should check my work, because i am tired and my algebra is sloppy.

OpenStudy (anonymous):

Ok thank you very much! :D